Given the demand for a product as qd=100-5p the supply is given as qs=-20+3p. You are told equilibrium is obtained at the point where qd=qs. The equilibrium quantity for product is?


1
Expert's answer
2021-06-03T18:35:05-0400

Solution:

At equilibrium: Quantity demanded (Qd) = Quantity supplied (Qs)

Quantity demanded = 100 – 5p

Quantity supplied = -20 + 3p

Therefore:

100 – 5p = -20 + 3p

100 + 20 = 3p + 5p

120 = 8p

P = "\frac{120}{8} = 15"


Equilibrium price (p) = 15


Substitute in the demand or supply function to derive equilibrium quantity:

100 – 5p

100 – 5(15) = 100 – 75 = 25


Equilibrium quantity of the product (q)= 25